OpenStudy (anonymous):

find slope at p find equation of tangent at point p y = 6 - x^2 p (2,2)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

confused about?

OpenStudy (anonymous):

this homework

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you mean \[6-x^{2}\] to be the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First, find \(\frac{dy}{dx}\).\[\frac{dy}{dx}=0+2x^{2-1}\] or \[\frac{dy}{dx}=2x\] This is because the derivative of a constant is 0 and the derivitive of a power is \[\frac{d}{dx}x^{n}=nx^{n-1}\] Now just solve at \(x=2\):\[\frac{dy}{dx}=2(2)=4\] This is the slope, and you already know the tanget passes through \((2,2)\). Using point-slope form: \[y-y_{0}=m(x-x_{0})\] \[y-2=4(x-2)\] or, in terms of y: \[y=4x-6\]